# compat=1.1.0

### 19 Dec 2023 v1

@@ V.2 Decoupling Limits
The couplings $h^{0}H^{+}H^{-}$ and $h^{0}h^{0}H^{+}H^{-}$ demonstrate sensitivity to distinct sets of Feynman diagrams. Specifically, $h^{0}H^{+}H^{-}$ exhibits sensitivity to T2, B2, B4, B6, V1, and V7, while $h^{0}h^{0}H^{+}H^{-}$ is responsive to the topologies V11 and V15 depicted in Figs. 1, 2 and 3. Examining Eq. (F) and Eq. (F), it is observed that the magnitudes of these couplings reach their maximum values when the parameter $\lambda_{4}$ is set to $1.82$, while their minimum values are obtained at $\lambda_{4}=-1.82$ in the carried out computations. Notably, the results presented in Figs. 67, and  8 suggest that the cross sections of $h^{0}h^{0}$ production experience enhancement when $\lambda_{4}>0$, despite the small value of $\lambda_{H^{0}h^{0}h^{0}}$ at the weak coupling decoupling limit. Additionally, the coupling $h^{0}h^{0}H^{++}H^{--}$ shows sensitivity to the Feynman diagrams V11 and V15 and Eq. (65) indicates this coupling maintains a constant and non-zero value throughout the calculations.
The couplings $h^{0}A^{0}A^{0}$ and $H^{0}A^{0}A^{0}$ exhibit sensitivity to specific Feynman diagrams, namely T1, T2, T3, V1, and V2, as illustrated in Fig. 1 and Fig. 3. According to Eq. (F), as $\lambda^{HTM}_{H^{0}A^{0}A^{0}}\approx\sqrt{2}\mu\tan^{2}{\beta}$ for small $\mu$, $\lambda^{HTM}_{H^{0}A^{0}A^{0}}$ also approaches the weak-coupling decoupling limit when $\tan{\beta}\ll 1$ and $\lambda_{2}+\lambda_{3}=1+(-1)=0$. For the values of $\mu\in[0.2,2]$, $2.8\times 10^{-5}<\lambda^{HTM}_{H^{0}A^{0}A^{0}}<2.8\times 10^{-3}$, but the $H^{0}A^{0}A^{0}$ coupling becomes substantial for large $\mu$. Despite the relatively small magnitude of $H^{0}A^{0}A^{0}$ in the calculations, Eq. (F) reveals that the magnitude of $h^{0}A^{0}A^{0}$ reaches its maximum when $\lambda_{4}$ is greater than 0. Conversely, this coupling attains its minimum values at $\lambda_{4}=-1.82$ in our calculations. Moving on to the couplings $A^{0}A^{0}H^{++}H^{--}$ and $A^{0}A^{0}H^{+}H^{-}$, their sensitivity lies in Feynman diagrams labeled as V11 and V15 in Fig. 3. According to Eq. (67), $\lambda_{A^{0}A^{0}H^{++}H^{--}}^{HTM}$ remains non-zero, while Eq. (F) indicates that $\lambda_{A^{0}A^{0}H^{+}H^{-}}^{HTM}\approx-3/2(\lambda_{1}+\lambda_{4})\tan^{2}{\beta}$ is relatively small. Notably, the results presented in Figs. 910 and  11 suggest an enhancement in the cross sections of $A^{0}A^{0}$ production when $\lambda_{4}$ is greater than 0.
@@
$\frac{\Delta\sigma}{\sigma_{SM}}=\Big{|}\frac{\sigma_{HTM}(\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0})}{\sigma_{SM}(\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0})}-1\Big{|}$ (44)
@@ VI Conclusions
In our analysis, we demonstrate that the cross sections for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ exhibit enhancement when $M_{H^{0}}\approx M_{A^{0}}>M_{H^{\pm}}>M_{H^{\pm\pm}}$, and the cross sections diminish to lower values when $M_{H^{\pm}}<M_{H^{\pm\pm}}<M_{H^{0}}\approx M_{A^{0}}$. To perform these calculations, the $\mathcal{P}_{1}$ parameter space was employed, setting $\alpha\approx 0$ to ensure that the light CP-even Higgs behaves similarly to the SM Higgs through three distinct approaches. Our first approach involved convoluting the partonic cross sections with the Weizsäcker-Williams approximation (LO), followed by the second approach, where we convoluted with the Weizsäcker-Williams approximation corrected up to the second order, $\mathcal{O}(\alpha_{e}^{2}t^{2})$. Lastly, we employed the EW PDF set and discussed the differences among the results generated by each approach. Throughout our comprehensive calculations, the behavior of $\lambda_{hhh}^{HTM}$ closely aligns with $\lambda_{hhh}^{SM}$ numerically. Here, we explicitly demonstrate the significant contributions of $h^{0}H^{+}H^{-}$, $h^{0}h^{0}H^{+}H^{-}$, and $h^{0}A^{0}A^{0}$ in enhancing the cross sections of $\phi\phi$ production mechanisms. We have observed that the magnitude of the $H^{0}h^{0}h^{0}$ coupling, where $\sigma_{HTM}\sim\sigma_{SM}$, remains fixed for each hierarchy. We explored this using the aforementioned PDFs as well. These findings provide crucial insights into the comparable numerical behavior of the HTM and the SM regarding the $\lambda_{hhh}$ parameter.
@@ Acknowledgements.
Appendix A Decay width of H
The total decay width of the heavy Higgs in the Higgs Triplet Model was calculated using the following relationships. The decay rates for the Higgs boson decaying into two gauge bosons, $VV$ $(V=W,Z)$, are given by:
$\Gamma(H\rightarrow VV)=\frac{\delta_{V}}{512\pi}\eta_{V}^{2}M_{H}^{3}\left[1-\frac{4M_{V}^{2}}{M_{H}^{2}}+\frac{12M^{2}_{Z}}{M_{H}^{4}}\right]\beta\left(\frac{M^{2}_{V}}{M_{H}^{2}}\right),$ (45)
where $\delta=2$ for W and $\delta=1$ for Z, and
@@ Appendix A Decay width of H
$\Gamma(H\rightarrow ff)=\frac{g^{2}}{32\pi M_{W}^{2}}M_{H}^{2}M_{f}^{2}N^{f}_{c}\beta\left(\frac{M_{f}^{2}}{M_{H}^{2}}\right)^{3}\sin^{2}{\alpha},$ (48)
where $M_{f}$ is the fermion mass and $N^{f}_{c}$ is the color factor. Note that,

### 05 Apr 2024 v2

@@ V.2 Decoupling limits
The couplings $h^{0}H^{+}H^{-}$ and $h^{0}h^{0}H^{+}H^{-}$ demonstrate sensitivity to distinct sets of Feynman diagrams. Specifically, $h^{0}H^{+}H^{-}$ exhibits sensitivity to T2, B2, B4, B6, V1, and V7, while $h^{0}h^{0}H^{+}H^{-}$ is responsive to the topologies V11 and V15 depicted in Figs. 1, 2, and 3. Examining Eq. (F) and Eq. (F), it is observed that the magnitudes of these couplings reach their maximum values when the parameter $\lambda_{4}$ is set to $1.82$, while their minimum values are obtained at $\lambda_{4}=-1.82$ in the carried out computations. Notably, the results presented in Figs. 67, and  8 suggest that the cross sections of $h^{0}h^{0}$ production experience enhancement when $\lambda_{4}>0$, despite the small value of $\lambda_{H^{0}h^{0}h^{0}}^{HTM}$ at the weak coupling decoupling limit. Additionally, the coupling $h^{0}h^{0}H^{++}H^{--}$ shows sensitivity to the Feynman diagrams V11 and V15 and Eq. (99) indicates this coupling maintains a constant and nonzero value throughout the calculations.
The couplings $h^{0}A^{0}A^{0}$ and $H^{0}A^{0}A^{0}$ exhibit sensitivity to specific Feynman diagrams, namely T1, T2, T3, V1, and V2, as illustrated in Fig. 1 and Fig. 3. According to Eq. (F), as $\lambda^{HTM}_{H^{0}A^{0}A^{0}}\approx\sqrt{2}\mu\tan^{2}{\beta}$ for small $\mu$, $\lambda^{HTM}_{H^{0}A^{0}A^{0}}$ also approaches the weak-coupling decoupling limit when $\tan{\beta}\ll 1$ and $\lambda_{2}+\lambda_{3}=1+(-1)=0$. For the values of $\mu\in[0.2,2]$, $2.8\times 10^{-5}<\lambda^{HTM}_{H^{0}A^{0}A^{0}}<2.8\times 10^{-3}$, but the $H^{0}A^{0}A^{0}$ coupling becomes substantial for large $\mu$. Despite the relatively small magnitude of $H^{0}A^{0}A^{0}$ in the calculations, Eq. (F) reveals that the magnitude of $h^{0}A^{0}A^{0}$ reaches its maximum when $\lambda_{4}$ is greater than 0. Conversely, this coupling attains its minimum values at $\lambda_{4}=-1.82$ in our calculations. Moving on to the couplings $A^{0}A^{0}H^{++}H^{--}$ and $A^{0}A^{0}H^{+}H^{-}$, their sensitivity lies in Feynman diagrams labeled as V11 and V15 in Fig. 3. According to Eq. (101), $\lambda_{A^{0}A^{0}H^{++}H^{--}}^{HTM}$ remains nonzero, while Eq. (F) indicates that $\lambda_{A^{0}A^{0}H^{+}H^{-}}^{HTM}\approx-3/2(\lambda_{1}+\lambda_{4})\tan^{2}{\beta}$ is relatively small. Notably, the results presented in Figs. 910, and  11 suggest an enhancement in the cross sections of $A^{0}A^{0}$ production when $\lambda_{4}$ is greater than 0.
@@
$\frac{\Delta\sigma}{\sigma_{SM}}=\frac{\sigma_{HTM}(\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0})}{\sigma_{SM}(\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0})}-1$ (44)
@@ VI Conclusions
In our analysis, we demonstrate that the cross sections for $\mu^{+}\mu^{-}\rightarrow\gamma\gamma\rightarrow h^{0}h^{0},A^{0}A^{0}$ exhibit enhancement when $M_{H^{0}}\approx M_{A^{0}}>M_{H^{\pm}}>M_{H^{\pm\pm}}$, and the cross sections diminish to lower values when $M_{H^{0}}\approx M_{A^{0}}<M_{H^{\pm}}<M_{H^{\pm\pm}}$. To perform these calculations, the $\mathcal{P}_{1}$ parameter space was employed, setting $\alpha\approx 0$ to ensure that the light CP-even Higgs behaves similarly to the SM Higgs through three distinct approaches. Our first approach involved convoluting the partonic cross sections with the Weizsäcker-Williams approximation (LO), followed by the second approach, where we convoluted with the Weizsäcker-Williams approximation corrected up to the second order, $\mathcal{O}(\alpha_{e}^{2}t^{2})$. Lastly, we employed the EW PDF set and discussed the differences among the results generated by each approach. Throughout our comprehensive calculations, the behavior of $\lambda_{hhh}^{HTM}$ closely aligns with $\lambda_{hhh}^{SM}$ numerically. Here, we explicitly demonstrate the significant contributions of $h^{0}H^{+}H^{-}$, $h^{0}h^{0}H^{+}H^{-}$, and $h^{0}A^{0}A^{0}$ in enhancing the cross sections of $\phi\phi$ production mechanisms. We have observed that the magnitude of the $H^{0}h^{0}h^{0}$ coupling, where $\sigma_{HTM}\sim\sigma_{SM}$, remains fixed for each hierarchy. We explored this using the aforementioned PDFs as well. These findings provide crucial insights into the comparable numerical behavior of the HTM and the SM regarding the $\lambda_{hhh}$ parameter.
@@ Acknowledgements.
Appendix A DECAY WIDTH OF $H^{0}$ AND $A^{0}$
In this appendix, we provide some useful information about the decay widths and branching ratios of $H^{0}$ and $A^{0}$ that could be essential for further discussions. Figs 15- 19 show the numerical results of the decay widths and branching ratios according to the parameter choices in Section V. The branching ratios of $H^{0}\rightarrow t\bar{t},W^{+}W^{-},Z^{0}Z^{0}$, and $A^{0}\rightarrow t\bar{t}$ were calculated using FeynRules [31].
The decay channels of $A^{0}/H^{0}\rightarrow\gamma\gamma,\gamma Z^{0},gg$ are given by analytic expressions 45 and 46. The contributions from $H^{\pm}$ and $H^{\pm\pm}$ should be considered to calculate the decay widths of $H^{0}\rightarrow\gamma\gamma,\gamma Z^{0}$ in the context of HTM.
$\Gamma(A^{0}/H^{0}\rightarrow\gamma\gamma,gg)=\frac{|\mathcal{M}(A^{0}/H^{0}\rightarrow\gamma\gamma,gg)|^{2}}{32\pi M_{A^{0}/H^{0}}},$ (45)
$\Gamma(A^{0}/H^{0}\rightarrow\gamma Z^{0})=\frac{|\mathcal{M}(A^{0}/H^{0}\rightarrow\gamma Z^{0})|^{2}}{16\pi M_{A^{0}/H^{0}}}\Big{(}1-\frac{M_{Z}^{2}}{M_{A^{0}/H^{0}}^{2}}\Big{)}.$ (46)
The analytic expression for $|\mathcal{M}|$ is given by,
$|\mathcal{M}(A^{0}/H^{0}\rightarrow XX)|^{2}$ $=|\mathcal{M}_{++}(A^{0}/H^{0}\rightarrow XX)|^{2}$
$\quad+|\mathcal{M}_{--}(A^{0}/H^{0}\rightarrow XX)|^{2},$ (47)
where $XX=\gamma\gamma,gg,\gamma Z^{0}$. According to FormCalc output, we were able to confirm that $\mathcal{M}_{+-}=\mathcal{M}_{-+}=0$ and $|\mathcal{M}_{+-}|^{2}=|\mathcal{M}_{-+}|^{2}$. The analytic expressions for $\mathcal{M}_{++}$ are shown in 55-68. In these expressions, $f$ and $q$ represent fermions and quarks, respectively, where $N_{f}^{c}=3$ for quarks and $1$ for leptons. $Q_{f}$ denotes the charge of the Standard Model fermions. The following scalar functions used in these expressions are evaluated with LoopTools, Ref. [34].
$B_{0}(p^{2},m_{1}^{2},m_{2}^{2})=\int\frac{d^{D}q}{i\pi^{2}}\frac{1}{(q^{2}-m_{1}^{2})((q+p)^{2}-m_{2}^{2})},$ (48)
$C_{0},C^{\mu},C^{\mu\nu}(p_{1}^{2},p_{2}^{2},(p_{1}+p_{2})^{2},m_{1}^{2},m_{2}^{2},m_{3}^{2})$ $=$
$\int\frac{d^{D}q}{i\pi}\frac{1,q^{\mu},q^{\mu}q^{\nu}}{(q^{2}-m_{1}^{2})((q+p_{1})^{2}-m_{2}^{2})((q+p_{1}+p_{2})^{2}-m_{3}^{2})}.$ (49)
The decay rates for the Higgs boson decaying into two gauge bosons, $VV$ $(V=W^{\pm},Z^{0})$, are given by, Ref. [39]:
$\Gamma(H^{0}\rightarrow VV)=\frac{\delta_{V}}{512\pi}\eta_{V}^{2}M_{H}^{3}\left[1-\frac{4M_{V}^{2}}{M_{H}^{2}}+\frac{12M^{2}_{Z}}{M_{H}^{4}}\right]\beta\left(\frac{M^{2}_{V}}{M_{H}^{2}}\right),$ (50)
where $\delta=2$ for $W^{\pm}$ and $\delta=1$ for $Z^{0}$, and
@@ Appendix A DECAY WIDTH OF $H^{0}$ AND $A^{0}$
$\Gamma(H^{0}\rightarrow f\bar{f})=\frac{g^{2}}{32\pi M_{W}^{2}}M_{H}^{2}M_{f}^{2}N^{c}_{f}\beta\left(\frac{M_{f}^{2}}{M_{H}^{2}}\right)^{3}\sin^{2}{\alpha},$ (53)
where $M_{f}$ is the fermion mass. Note that,